OpenStudy (dani_rose):

I NEED HELP!! What is 2+2=?

OpenStudy (dani_rose):

@bigpapa1

OpenStudy (anonymous):

4

OpenStudy (dani_rose):

Impossible....

OpenStudy (anonymous):

lol

OpenStudy (bigpapa1):

OMG it's 4 lol @Dani_Rose

OpenStudy (dani_rose):

Don't judge me lol

OpenStudy (daniee_bruhh):

Guys don't over think it. It's clearly 7 XD

OpenStudy (anonymous):

lol

OpenStudy (bigpapa1):

Since when did you not know 2+2? @Dani_Rose

OpenStudy (anonymous):

lol

OpenStudy (daniee_bruhh):

Her whole life lol

OpenStudy (dani_rose):

Wow..... @Daniee_Bruhh

OpenStudy (daniee_bruhh):

EVERYTHING I KNOW IS A LIE IF 2+2=4

OpenStudy (daniee_bruhh):

I'm a genius huh

OpenStudy (dani_rose):

-hits report abuse button- lol

OpenStudy (daniee_bruhh):

Ready to be fanned in 3...2...1 lol

OpenStudy (daniee_bruhh):

lol

OpenStudy (dani_rose):

My first question EVER ;-;

OpenStudy (dani_rose):

lol i just wanna say thank you for being apart of this big step in my life

OpenStudy (anonymous):

NO. 2+2=3

OpenStudy (bigpapa1):

No problem lol

OpenStudy (daniee_bruhh):

@bigpapa1 we're honered (I can't spell ) Honored XD

OpenStudy (dani_rose):

@Getalampbro Got it right

OpenStudy (daniee_bruhh):

I thought I got it right? lol

OpenStudy (dani_rose):

Challenge: Show the steps

OpenStudy (dani_rose):

no you didn't @Daniee_Bruhh

OpenStudy (daniee_bruhh):

\[easy : 1+1=2 so 2+2=7 \]

OpenStudy (anonymous):

2+2=3. Because reasons

OpenStudy (dani_rose):

"Because Reasons" lol CORRECT

OpenStudy (daniee_bruhh):

lol I actually showed the work so I'm right

OpenStudy (dani_rose):

Well your work was wrong

OpenStudy (daniee_bruhh):

WOW *logs off*

OpenStudy (dani_rose):

Hater lol jk

OpenStudy (anonymous):

hah

OpenStudy (anonymous):

burned

OpenStudy (bigpapa1):

The fact that you guys are siblings is really showing atm

OpenStudy (dani_rose):

So if 2+2=3... Then what is 0+0=?

OpenStudy (bigpapa1):

0

OpenStudy (anonymous):

retard

OpenStudy (dani_rose):

Wrongggggg

OpenStudy (bigpapa1):

retard is a rude word @jerlin1103

OpenStudy (daniee_bruhh):

bro I'm so confused right now no joke lol I have no idea what is happening

OpenStudy (dani_rose):

o=0

OpenStudy (daniee_bruhh):

WAIT WHO'S A RETARD?

OpenStudy (dani_rose):

No mean language

OpenStudy (dani_rose):

Or Get banned

OpenStudy (bigpapa1):

I don't know but someone just said it it wasn't me

OpenStudy (dani_rose):

lol ik

OpenStudy (anonymous):

lol sorry sir

OpenStudy (daniee_bruhh):

lol I'll be the bigger bear and be the retard XD

OpenStudy (dani_rose):

Sir lol im a panda

OpenStudy (bigpapa1):

Why use the word sir?

OpenStudy (daniee_bruhh):

that's a great question.... hmmmm

OpenStudy (daniee_bruhh):

@bigpapa1 I have a question that's gonna sound dumb lol can you help?

OpenStudy (anonymous):

idk

OpenStudy (bigpapa1):

Yes I probably can @Daniee_Bruhh

OpenStudy (dani_rose):

lol probably

OpenStudy (anonymous):

lol ok i want to see what it is

OpenStudy (daniee_bruhh):

what does the (16) mean next to my profile picture?

OpenStudy (dani_rose):

Smart Score

OpenStudy (bigpapa1):

it's your smart score

OpenStudy (dani_rose):

Beat you to it

OpenStudy (anonymous):

lol

OpenStudy (daniee_bruhh):

lol so i'm only 16% smart? While you guys are mini Einsteins lol

OpenStudy (dani_rose):

Shower meh wif chur medallllss

OpenStudy (dani_rose):

lol im baby Einstein...

OpenStudy (daniee_bruhh):

lol I read that wrong. I saw shower me with your cheeseballs (don't ask why I thought that cuz I have no idea)

OpenStudy (anonymous):

I am waiting for a person to tell ous to close this question

OpenStudy (daniee_bruhh):

they do that?

OpenStudy (bigpapa1):

Yes they do it's happened to me a couple times

OpenStudy (anonymous):

Yea the helpers

OpenStudy (daniee_bruhh):

lol they sound like evil people in books... "Yes they call them the Helpers."

OpenStudy (anonymous):

They helped me once i told them to f off

OpenStudy (daniee_bruhh):

why if they helped you?

OpenStudy (bigpapa1):

OMG

OpenStudy (daniee_bruhh):

lol did you say OMG because you know that they're coming?

OpenStudy (dani_rose):

it shall neva be closed

OpenStudy (anonymous):

I said that Because yolo

OpenStudy (bigpapa1):

lol #YOLO

OpenStudy (daniee_bruhh):

oh lol okay "makes sense"

OpenStudy (dani_rose):

If they come and say to shut it down im gonna say its @bigpapa1 Question not mine lol

OpenStudy (bigpapa1):

Bruh I've already done this once :P

OpenStudy (daniee_bruhh):

lol I'm starting to be afraid of this site

OpenStudy (dani_rose):

lol then you have experience

OpenStudy (anonymous):

LOL MY BROTHER SAID STFU 2+2=4

OpenStudy (dani_rose):

He's wrong

OpenStudy (anonymous):

My brother is dumb

OpenStudy (daniee_bruhh):

LOL don't listen to him... He knows nothing XD

OpenStudy (dani_rose):

He must be infected with Smartkiditis...

OpenStudy (dani_rose):

it's a rare disease

OpenStudy (anonymous):

The worst dieses

OpenStudy (dani_rose):

the worst indeed

OpenStudy (anonymous):

He said you guys must be retarded

OpenStudy (dani_rose):

lol i am

OpenStudy (bigpapa1):

he sounds mean

OpenStudy (dani_rose):

i just thought of poptarts..

OpenStudy (anonymous):

Yep that my life